OpenStudy (anonymous):

What equation results from completing the square and then factoring? x2 + 10x = 15

OpenStudy (imstuck):

As long as the coefficient on the x^2 term is a 1, you can complete the square. Take half the x term 10, which is 5, square the 5 to get 25, and then add the 25 to both sides. This results in an equation that looks like this, before simplifying:

OpenStudy (vispla22):

are you sure its not 25 at the end?

OpenStudy (imstuck):

\[x ^{2}+10x+25=15+25\]

OpenStudy (imstuck):

Then, simplifying:

OpenStudy (imstuck):

\[(x+5)^{2}=40\]